At the north campus of a small liberal arts college, 10% of the students are women. At the south campus, 50% of the students are women. The campuses are merged into one east campus. If 40% of the 1200 students at the east campus are women, how many students did the north and south campuses have before the merge?
Answer by nyc_function(2741)

x + y = 1200
y = 1200 - x
x= 1200 - y
.1x + .5y = 480
.1x + .5(1200 - x) = 480
.1x + 600 - .5x = 480
-.4x + 600 = 480
-.4x = -120
x = 300
therefore.. 10% of women on the north campus equal 300 students..
to find the whole amount divides.. 300/.1 = 3000 students on the north campus before the divide
.1x + .5y = 480
.1(1200 - y) + .5y = 480
120 - .1y + .5y = 480
120 + .4y = 480
.4y = 360
y = 900
therefore 50% of students before the merge that were women was 900 students.. to find the whole amount of students on the east side divide.. 900/.5 = 1800 students on the north side before the merge.